The perimeter of an isosceles triangle is 28cm. What is the maximum area possible? 


1
Expert's answer
2021-10-04T19:31:54-0400

Let "x" denotes the length of a leg of a triangle. Since the perimeter is 28 and the triangle is isoscales, the length of the base is "28-x-x=28-2x." Let "h" be the length of the altitude to the base. Since the triangle is isoscales, the altitude and the median to the base are coinside. By Pythagorean theorem, "h=\sqrt{x^2-(\frac{28-2x}2)^2}=\sqrt{x^2-(14-x)^2}=\sqrt{x^2-196 +28x-x^2}=\sqrt{28x-196}=2\sqrt{7x-49}."

Then the area of triangle is "A(x)=\frac{1}2(28-2x)2\sqrt{7x-49}=2(14-x)\sqrt{7x-49}."

It follows that "A'(x)=2(-\sqrt{7x-49}+(14-x)\frac{7}{2\sqrt{7x-49}})
=2\frac{-2(7x-49)+7(14-x)}{2\sqrt{7x-49}}
=\frac{196-21x}{\sqrt{7x-49}}."

If "A'(x)=0" then "196-21x=0," and hence "x=\frac{196}{21}=\frac{28}3." Since "A'(x)>0" for "x<\frac{28}3" and "A'(x)<0" for "x>\frac{28}3," we conclude that "x=\frac{28}3" is the point of maximum.

Taking into account that "A(\frac{28}3)=2(14-\frac{28}3)\sqrt{7\frac{28}3-49}=\frac{28}3\sqrt{\frac{49}3}=\frac{196\sqrt{3}}{9}," we conclude that "\frac{196\sqrt{3}}{9}" is the maximum area possible.